If a boy on a motorbike travels to school at 20km/hr reaches the school 13mins earlier than his sister who is on a bicycle going at 12km/hr, how far away is the school?
This information was given to me by my kid and he says that this was all the information given. I am unsure whether there is more info or not. If you cannot solve it due to lack of information can you please send me an email saying there is not enough info. This question has really been bugging me for some days now.

Answer by mananth(16949)

distance to school = x km
Motorbike time =x/20
Bicycle time =x/12
Difference in time =13/60 =0.22 hour
Bicycle time - motorbike time = 0.22
x/12-x/20=0.22

common denominator = 240
multiply by LCD
20x-12x=0.22*240
8x=52.8
/8
x=6.6 km